OpenStudy (anonymous):

Solving quadratic equations with complex solutions. I dont understand this! Solve x2 - 7x = -13.

OpenStudy (anonymous):

you have to use the formula y= \[y= x +or - \sqrt{b^2-4ac}\] all over 2a

OpenStudy (anonymous):

A. x= \[-7\pm i \sqrt{6} \over 2\] B. x= -7 \[\pm i \sqrt{3} \over 2\] C. x= \[7\pm i \sqrt{6} \over 2\] D. x= \[7 \pm i \sqrt{3} \over 2\]

OpenStudy (anonymous):

Could you explain how to do that

OpenStudy (anonymous):

I dont get much help from my teacher..

OpenStudy (anonymous):

@Liverbean